# Distilling Large Language Models for Network Active Queue Management

### v1

@@ II-A Architecture of AQM-LLM
5:$conv_{1}$, …, $conv_{n}$ as Conv1d(1, feature_dim, conv_size) for sequence data
6:$embed\_feature$, …, $embed\_feature\_n$ as     Linear(feature_dim, plm_embed_size) where     plm_embed_size depends on LLM token size.
@@
Feature Encoder: To extract features from raw input data, we employ specialized feature encoders that are tailored to each type of data. A key design choice is leveraging pre-existing, optimized encoders rather than manually building them from scratch. In this implementation, as outlined in Algorithm 1, we adopt the following approach: i) Scalar Data. Fully connected layers are used for processing scalar inputs, as shown in Algorithm 1, line 4. These layers efficiently handle single-value metrics such as queue size or drop probability; ii) Time-Series and Sequence Data. For sequential data, such as queue delay trends, we use 1-dimensional convolutional neural networks (CNNs) to capture temporal patterns, as depicted in line 5 of Algorithm 1. This modular design ensures efficient feature extraction across diverse input modalities while maintaining compatibility with the LLM framework.
Linear Projection: Although the feature encoders are highly capable, the dimensions of the extracted features may not match the token space required by the LLM. For example, Llama2 expects an input dimension of 4096 [9]. To resolve this mismatch, we utilize trainable linear or embedding layers (as shown in Algorithm 1, line 16) to project the extracted features into the appropriate token space. These embedding layers learn to map features to token-like embeddings compatible with the LLM, enabling seamless integration. Additionally, we apply layer normalization [22] to the output embeddings, ensuring improved training stability and performance. This approach ensures that the data is optimally formatted for the LLM while maintaining robust and efficient processing.
II-B L4S Language Modelling Head
Figure 4: L4S-LM Head
@@ II-B L4S Language Modelling Head
1:Input: $input\_embeddings,attention\_mask,$
2:$action\_embed\_pos,residual\_flag$
@@
4:Initialize L4S LM head:
5:action_head $\leftarrow$ Linear($plm\_embed\_size,actionCount$)
@@
Using the above state encoder, the LLM is able to transform the raw input data of different modalities into embeddings or high-level features. These features are then directed to the L4S-LM head to directly generate answers. We designed a custom L4S Language Modelling head, implemented as a trainable linear layer to generate task-specific answers based on the processed features from the LLM. Unlike the LM head, our custom L4S-LM head constrains answers to fall within a valid range of possible answers, ensuring the reliability of the LLM in our task. For example, in our case, we define three discrete actions: 0 refers to enqueuing the packet, 1 refers to dropping a packet from the queue, and 2 indicates applying the CE(1) flag to signal congestion for L4S enqueue management.
Traditionally, the LM head requires multiple rounds of inference to generate a single valid response, which leads to significant delays in answer generation. To address this challenge, we developed the L4S-LM head (refer to line 5 in Algorithm 2), to generate a valid response in a single inference round by predicting a probability distribution over possible actions. This new approach is inspired by speculative decoding, where instead of generating responses token by token, the model predicts multiple possible outcomes simultaneously, significantly enhancing efficiency without compromising response validity. The advantage of this method lies in its ability to provide quick and accurate responses, in network congestion control and AQM, where low latency is extremely crucial for optimal system performance. Longer inference times would defeat the purpose of utilizing LLMs in the first place. Figure 4 illustrates the traditional LM head and the L4S-LM head for comparison.
II-C Low-Rank Adaptation
In this section, we propose a data-driven low-rank L4S adaptation to efficiently fine-tune our LLM to reduce computation overhead and answer generation time for our AQM-specific task. This adaptation allows us to use an offline approach to fine-tune our model, instead of having to interact with the network environment and increase overhead directly. It comprises two main components (A) a data-driven pipeline for AQM decision-making and (B) a low-rank data adaptation approach (LoRA) which reduces the number of parameters that need to be fine-tuned from 7B parameters to just 1% of the parameters, for faster fine-tuning. LoRA has two advantages (A) it decreases the time for fine-tuning significantly, and (B) It helps us retain the foundational knowledge of LLM.
II-C1 RL-based L4S Adaptation
In the case of a task similar to the typical Markov Decision Process, instead of using the typical traditional RL approach, we adopt a data-driven RL technique. The standard RL technique is impractical due to the significant interaction time required between the LLM and the environment. To overcome this, we propose a modified RL adaptation pipeline based on the efficient data-driven RL technique [23, 24, 25]. This method eliminates the need for any real-time interaction by utilizing pre-collected experience datasets from existing networking algorithms. This approach enables fine-tuning the model to optimize reward. The intuitive rationale behind this approach is to allow the LLM to improve its policies by observing the behaviours of existing networking algorithms [24]. It learns why certain actions lead to better performance, as well as understanding why certain actions result in poor performance [25]. In this method, we only collect all our data at once to use throughout our training process, significantly decreasing the training time for our LLM.
The proposed RL adaptation pipeline functions by first collecting experience trajectories from an existing policy (DualPI2 from FreeBSD AQM) into a dataset $D_{\text{rl}}=\{\tau_{1},\dots,\tau_{|D_{\text{rl}}|}\}$, where each trajectory $\tau=\{r_{t},s_{t},a_{t}\}_{t=1}^{T}$ is composed of rewards $r$, states $s$ and actions $a$, where $T$ denotes the episode length. The actions are defined as, 0 - Enqueue, 1 - Drop, and 2 - MARKECN. The state consists of 8 features: {queue_type, burst_allowance, drop_probability, current_queue_delay, accumulated_probability, total_bytes , total_drops, packet_length}.
TABLE II: State Features
Features Definition
queue_type defines type of queue - Classic (0) or L4S (1)
burst_allowance Maximum allowed burst size of packets before congestion control measures are applied.
drop_prob Probability of dropping a packet when the queue is congested.
current_qdelay Current queue delay, which is the time a packet spends in the queue before being transmitted.
accu_prob Accumulated probability value used in FQ-PIE to determine the drop probability for incoming packets.
len_bytes Average length (in bytes) of the queue during the measurement interval.
drops Number of packets dropped during the measurement interval.
pkt_len length of enqueued packet
[…]
In the trajectory, the reward $r_{t}$ is substituted by the return $R_{t}=\sum_{i=t}^{T}r_{i}$, representing the cumulative rewards generated from the state $s_{t}$ moving forward. Furthermore, in certain tasks, the state or action may comprise multiple components of information. For example, in the context of L4S, the state includes historical queue delays. This complexity requires a finer level of discretization of both state and action spaces to ensure that all relevant aspects of the system are adequately captured.: $s_{t}=\{s_{t}^{1},\dots,s_{t}^{n}\},a_{t}=\{a_{t}^{1},\dots,a_{t}^{m}\}$. The updated representation of the trajectory is then expressed as:
$\tau=\{R_{t},s_{t}^{1},\dots,s_{t}^{n},a_{t}^{1},\dots,a_{t}^{m}\}_{t=1}^{T}$ (4)
Following the trajectory representation presented earlier, we fine-tune the LLM to learn the distribution of cumulated returns. For each training step, we randomly sample batches of data from the dataset:
$d=\{R_{t},s_{t}^{1},\dots,s_{t}^{n},a_{t}^{1},\dots,a_{t}^{m}\}_{t=t-w+1}^{T}$ (5)
Here, $w$ denotes the context window used to learn the return distribution. The LLM is then provided with data $d$ to generate actions $\{\hat{a}_{t}^{1},\dots,\hat{a}_{t}^{m}\}_{t=t-w+1}^{T}$. In this scenario, each component of return, action and state is treated as distinct modalities and processed independently. Then, the training loss can be calculated as follows:
$L_{rl}=\frac{1}{w}\sum_{t=1}^{w}\sum_{j=1}^{m}F_{rl}(a_{t}^{j},\hat{a}_{t}^{j})$ (6)
The function $F_{rl}$ measures the discrepancy between action $a_{i}^{j}$ and the generated action $\hat{a}_{i}^{j}$. For this, we utilize either cross-entropy (CE) for categorical actions or mean squared error (MSE) for continuous action values, , akin to regression tasks.
The training aims to model the distribution of returns conditioned on specific states, thus generating action sequences that achieve the desired returns [26]. During the inference stage, to guide the LLM towards the desired performance, a target return is specified.
II-C2 Low-rank L4S adaptation
Figure 5: Data-driven Low-Rank L4S adaptation (proposed LoRA) Diagram
We use LoRA to decrease training time by reducing the number of trainable parameters using a technique called Parameter-Efficient Tuning (Peft), in which we freeze the parameters of the LLM and incorporate additional low-rank matrices to approximate the required adjustments for domain-specific adaptation. We will denote the pre-trained parameters of the LLM as $\Phi_{0}$, and $\Delta\Phi$ as the trainable parameters, which we update and optimize to get the final model parameters as $\Phi=\Phi_{0}+\Delta\Phi$. Thus, for each pre-trained matrix $W_{0}$ in $\Phi_{0}$ of dimension $d\times k$, we hypothesise a low rank $r\ll\min\{d,k\}$ and construct two low-rank matrices $A$ and $B$ of dimensions $d\times r$ and $r\times k$, respectively, to approximate the update as $W=W_{0}+\Delta W=W_{0}+AB$.
@@ II-C2 Low-rank L4S adaptation
We freeze the pre-trained matrix $W_{0}$ during adaptation and confine all the updates to matrices $A$ and $B$ as shown in Algorithm 3. This strategy significantly lowers fine-tuning costs by reducing the demand on computational resources (such as GPU VRAM, CPU usage, and RAM) and shortening the overall training time. Freezing the pre-trained parameters preserves the existing pre-trained knowledge already present in the LLM. This enables the LLM to function as a robust foundational model, facilitating efficient adaptation to acquire a wide range of domain-specific knowledge.
@@ III-A Experimental Testbed Setup
We collect the experiences required for training our L4S-LLM by developing a testbed consisting of FreeBSD-based virtual machines, which are built and configured using VirtualBox as the virtualization platform. The Virtual Machines run FreeBSD version 14.1 for their Operating System. To create a network setup, we clone four virtual machines: two clients (Client 1 and Client 2), one router, and one server. Client 1 utilizes two different congestion control algorithms - Cubic and NewReno simultaneously, while Client 2 uses ECT(1)-enabled DCTCP congestion control algorithm. The router is configured with a custom-built DualPI2 algorithm that supports the L4S architecture, as described in the paper [5]. All the clients send data to the server through the router as illustrated in Figure 6. All the scripts and additional information required to set up the environment are made public in L4S-LLM repository.
@@
In our experiment to collect data, we configure the router with a bandwidth of 8 Mbps and a delay of 10 ms. We initiate two ‘iperf3’ connections for each TCP congestion control algorithm used in Client 1, with each connection utilizing two parallel flows via the ‘-P 2’ option in ‘iperf3’. Client 2 only utilizes a single flow for its ‘iperf3‘ connection. Each ‘iperf3‘ test runs for one minute. We repeat the experiment ten times and collect all the data. We have enabled ECN on all devices.
@@
To gather the data, we use a debugging tool that is directly embedded in the DualPI2 kernel code to store data from the kernel into the debug kernel logs. The relevant data is then extracted from these logs. If the size of the collected data exceeds 1000 KB, the log files will be truncated and rotated to prevent excessive growth. To increase the file size limits, we modify the configuration files for the ‘newsyslog’ service, found in /etc/newsyslog.conf.
@@
Once the data collection process is complete, we process the collected data into an experience pool to be fed into the LLM; see L4S-LLM experience pool [^5] https://github.com/MPTCP-FreeBSD/LLM_Gen_Exp_Pool implementation for details. We experiment with multiple foundation LLM models, including Llama2666https://www.llama.com/llama2[9] (7b), OPT (1.3b), GPT2777https://openai.com/index/better-language-models/[27] (125m) and T5-LLM [^8] https://github.com/google-research/text-to-text-transfer-transformer[28], which are fine-tuned for our L4S-LLM task. Among these, Llama2 demonstrated the best performance (see Sec III B) and was selected as the foundational LLM for extensive evaluation in simulations, where L4S-LLM is compared against the legacy AQM algorithms. To execute L4S-LLM, we utilize ‘Runpod’, an online cloud computing platform that provides access to powerful GPUs and scalable storage, ensuring efficient training and testing of our framework. We used two NVIDIA A40 GPUs, each equipped with 48 GB of VRAM, carried out on an Ubuntu 22.04 instance with the Cuda 11.8 version.
III-A1 Evaluation Metrics
We log key metrics during the training and testing of L4S-LLM, including loss, accuracy, reward, GPU/CPU utilization, computation time, RAM usage, and total training/testing time. To evaluate the model’s efficiency, we analyze loss and accuracy, which are calculated using cross-entropy loss for this multi-class classification problem. Cross-entropy loss measures the discrepancy between predicted probabilities and true labels, outputting a scalar value where lower values indicate better performance.
@@ III-A1 Evaluation Metrics
The training process employs mini-batches with gradient accumulation to enhance efficiency and stability. Gradient clipping is utilized to prevent gradient explosion by capping gradients to a predefined threshold during backpropagation. The model predicts a probability distribution over actions, represented by action.preds, and the predicted action is obtained using the argmax() function. True actions, extracted and normalized from the experience pool, are aligned with the predicted actions using a custom vectorization and conversion function to ensure compatibility for accuracy computation as follows:
@@ III-B1 Evaluation of Loss Dynamics in L4S-LLM
Figure 7 presents the evolution of the training loss trajectory of Llama2 (7b), OPT (1.3b), GPT2 (125m) and T5-LLM (220m), spanning 80 epochs. We used LoRA to reduce the parameter sizes of Llama2 and OPT, as both models have exceptionally large parameter sizes, necessitating longer training times. Llama2 demonstrates the most consistent and rapid convergence, achieving the lowest mean loss amongst other models as soon as epoch 5 and keeps on improving. OPT shows a gradual improvement in the mean loss and converges to a slightly higher mean loss, indicating a slower optimization. T5-LLM follows a similar trend to OPT, showing a gradual improvement but achieves a slightly lower loss than OPT and GPT2. GPT2 initially starts with the highest mean loss but rapidly converges to and follows a similar trend to OPT from epoch 40, and also converges to a similar final mean loss of OPT.
These results showcase L4S-LLM’s ability to adapt to complex network queue management strategies regardless of the LLM model being used. Using historical real-time metrics captured directly from the kernel, L4S-LLM can optimize and generalise well given the dynamic network conditions. Furthermore, Llama2 with its 7 billion parameter set achieves the lowest mean loss, while OPT and GPT2 come in second and T5-LLM seems to have the highest loss by small margins.

### v3

@@ II-D Architecture of AQM-LLM
[…]
6:embed_featureembed\_feature, …, embed_feature_nembed\_feature\_n as   Linear(feature_dim, plm_embed_size) where   plm_embed_size depends on LLM token size.
@@
II-D1 L4S-LLM State Encoder
Figure 5: L4S-LLM Encoder: Detailed architecture showing how different types of network metrics are processed through specialized encoding paths.
The L4S-LLM State Encoder transforms raw network metrics into a format compatible with transformer models. As shown in Figure 5, it consists of two main components:
  1. 1.
    Feature Encoder: Processes different types of network metrics through specialized paths
  2. 2.
    Linear Projector: Maps extracted features to the token space required by the LLM
Algorithm 1 details the encoding process. For scalar metrics like queue type or drop probability, fully connected layers (line 4, Algorithm 1) extract meaningful representations. For time-series data like queue delay trends, 1D CNNs (line 5, Algorithm 1) capture temporal patterns with kernel sizes $k\in\{3,5,7\}$ to extract multi-scale features.
The extracted features are then projected into the LLM’s token space using trainable embedding layers (line 16, Algorithm 1). For example, with Llama2, features are mapped to 4096-dimensional vectors compatible with the model’s input requirements.
II-D2 L4S Language Modelling Head
Figure 6: L4S-LLM Head
Traditional language model heads generate responses token-by-token, requiring multiple inference rounds and introducing unacceptable latency for AQM applications. Our custom L4S-LLM head as shown in Figure 6 addresses this by generating complete actions in a single inference step. See details in Algorithm 2.
@@ II-D2 L4S Language Modelling Head
1:Input: input_embeddings,attention_mask,input\_embeddings,attention\_mask,
2:action_embed_pos,residual_flagaction\_embed\_pos,residual\_flag
@@
4:Initialize L4SLLM head:
5:action_head $\leftarrow$ Linear(plm_embed_size,actionCountplm\_embed\_size,actionCount)
@@
The L4S-LLM head (line 5, Algorithm 2) is a trainable linear layer that maps transformer outputs directly to a probability distribution over the three possible actions: enqueue (0), drop (1), or ECN mark (2).666 This approach, inspired by speculative decoding, eliminates the need for token-by-token generation, reducing inference time from hundreds of milliseconds to under 40ms.
II-E Data-Driven Reinforcement Learning (RL) Pipeline
Recall that we employ RL-based adaptation pipeline which begins by collecting experience trajectories from the FreeBSD DualPI2 implementation. Each trajectory $\tau=\{r_{t},s_{t},a_{t}\}_{t=1}^{T}$ consists of states $s_{t}$ as 8-dimensional vectors of network metrics (including queue type, burst allowance, drop probability, current queue delay, accumulated probability, queue length in bytes, packet drops, and packet length), actions $a_{t}$ as discrete values (0: enqueue, 1: drop, 2: ECN mark), and rewards $r_{t}$ (for instance computed as $\text{pkt\_len}/(\text{qdelay}_{t}+1)$ in our implementation) to balance throughput and latency. For each trajectory, we evaluate the return $R_{t}=\sum_{i=t}^{T}\gamma^{i-t}r_{i}$ with discount factor $\gamma=0.95$, representing cumulative future rewards from state $s_{t}$, transforming the trajectory into $\tau=\{R_{t},s_{t}^{1},\ldots,s_{t}^{8},a_{t}\}_{t=1}^{T}$.
During training, we sample batches of data from the experience pool as $d=\{R_{t},s_{t}^{1},\ldots,s_{t}^{8},a_{t}\}_{t=t-w+1}^{T}$ where $w$ is the context window, empirically determined to balance historical information against computational efficiency. The LLM processes this data to generate predicted actions $\{\hat{a}_{t}\}_{t=t-w+1}^{T}$, with training loss calculated as $L_{rl}=\frac{1}{w}\sum_{t=1}^{w}F_{rl}(a_{t},\hat{a}_{t})$ where $F_{rl}$ is the cross-entropy loss function for our categorical action space. The model processes encoded states through the transformer architecture and generates action probabilities via the L4S-specific classification head, enabling the system to learn optimal congestion control policies from historical network behavior patterns.
II-F Low-Rank Adaptation (LoRA)
Figure 7: Low-Rank L4S adaptation (LoRA) Design: Visualization of how LoRA matrices are integrated with frozen pre-trained weights to enable efficient fine-tuning.
Fine-tuning large language models with billions of parameters is computationally prohibitive for networking applications. We address this using Low-Rank Adaptation (LoRA), which dramatically reduces the number of trainable parameters while preserving model quality. Figure 7 shows the LoRA fine-tuning architecture where low-rank matrices (A and B) integrate with frozen pre-trained weights, enabling parameter-efficient L4S-LLM adaptation to network-specific patterns without full model retraining. Algorithm 3 details the LoRA training process. For each pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ in the LLM, we introduce low-rank matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ with rank $r\ll\min(d,k)$. The effective weight matrix
$W=W_{0}+AB$ (3)
By freezing $W_{0}$ and only training $A$ and $B$, we reduce the number of trainable parameters from billions to millions—a 99% reduction for Llama2-7B.
@@ II-F Low-Rank Adaptation (LoRA)
For Llama2-7B, we use rank $r=128$, reducing trainable parameters from 7 billion to approximately 70 million while maintaining 97.56% accuracy. This approach preserves the pre-trained knowledge in the LLM while enabling efficient adaptation to the AQM domain.
@@ III-A Experimental Testbed Setup
We collect the experiences required for training our L4S-LLM by developing a testbed as shown in Figure 8 consisting of FreeBSD-based virtual machines, built and configured using VirtualBox and VMware virtualization platforms. The virtual machines run FreeBSD version 14.1 as their operating system. To create a network setup in VirtualBox, we clone four virtual machines: two clients (Client 1 and Client 2), one router, and one server. Client 1 utilizes two different congestion control algorithms - Cubic and NewReno simultaneously, while Client 2 uses ECT(1)-enabled DCTCP congestion control algorithm. We replace DCTCP with UDP Prague and repeat all the experiments in VMware. [^7] VirtualBox provides open-source accessibility, simpler network configuration options, better FreeBSD guest support through built-in additions, and lower resource requirements suitable for basic protocol testing. In contrast, VMware delivers superior network performance with enhanced throughput and lower latency, advanced virtual networking capabilities (distributed switches, VLAN trunking, traffic shaping), better hardware virtualization, and more reliable snapshot functionality for experimental state preservation. For performance-sensitive experiments like congestion control or AQM evaluation, VMware’s network stack more accurately emulates physical hardware behavior, while VirtualBox remains adequate for functional protocol testing with reduced resource overhead. The router is configured with a custom-built DualPI2 algorithm that supports the L4S architecture, as described in [5]. All clients send data to the server through the router as illustrated in Figure 8. All scripts and additional information required to set up the environment are made public in the L4S-LLM repository.
@@
In our experiment to collect data, we configure the router with a bandwidth of 8 Mbps and a delay of 10 ms. We initiate two iperf3 connections for each congestion control algorithm used in Client 1, with each connection utilizing two parallel flows via the -P 2 option in iperf3. Client 2 only utilizes a single flow for its iperf3 connection. Each iperf3 test runs for one minute. We repeat the experiment ten times and collect all the data. We have enabled ECN on all devices.
@@
To gather the data, we use a debugging tool that is directly embedded in the DualPI2 kernel code to store data from the kernel into the debug kernel logs. The relevant data is then extracted from these logs. If the size of the collected data exceeds 1000 KB, the log files will be truncated and rotated to prevent excessive growth. To increase the file size limits, we modify the configuration files for the newsyslog service, found in /etc/newsyslog.conf.
@@
Once the data collection process is complete, we process the collected data into an experience pool to be fed into the LLM; see L4S-LLM experience pool [^8] https://github.com/MPTCP-FreeBSD/LLM_Gen_Exp_Pool implementation for details. We experiment with multiple foundation LLM models, including Llama2999https://www.llama.com/llama2[9] (7b), OPT (1.3b), GPT2101010https://openai.com/index/better-language-models/[35] (125m) and T5-LLM [^11] https://github.com/google-research/text-to-text-transfer-transformer[36], which are fine-tuned for our L4S-LLM task. Among these, Llama2 demonstrated the best performance (see Section III-B) and was selected as the foundational LLM for extensive evaluation in simulations, where L4S-LLM is compared against the legacy AQM algorithms. To execute L4S-LLM, we utilize Runpod, an online cloud computing platform that provides access to powerful GPUs and scalable storage, ensuring efficient training and testing of our framework. We used two NVIDIA A40 GPUs, each equipped with 48 GB of VRAM, carried out on an Ubuntu 22.04 instance with CUDA 11.8 version.
@@
We log key metrics during L4S-LLM training and testing, including loss, accuracy, reward, GPU/CPU utilization, computation time, RAM usage, and total training/testing time. To evaluate the model’s efficiency, we analyze loss and accuracy, which are calculated using cross-entropy loss for this multi-class classification problem. Cross-entropy loss measures the discrepancy between predicted probabilities and true labels, outputting a scalar value where lower values indicate better performance. The training process employs mini-batches with gradient accumulation to enhance efficiency and stability. Gradient clipping is utilized to prevent gradient explosion by capping gradients to a predefined threshold during backpropagation. The model predicts a probability distribution over actions, represented by action.preds, and the predicted action is obtained using the argmax() function. True actions, extracted and normalized from the experience pool, are aligned with the predicted actions using a custom vectorization and conversion function to ensure compatibility for computing accuracy
@@
To ensure robust generalization beyond training scenarios, we implemented multiple strategies to counteract limitations of supervised distillation. We systematically varied key network parameters (bandwidth, RTT, queue sizes, traffic patterns) during data collection to prevent model memorization. Our approach incorporated queue perturbations, temporal shifting of metrics, and diverse traffic traces from multiple congestion control algorithms to create novel training examples.
@@ III-B1 Evaluation of Loss Dynamics in L4S-LLM
Figure 9 demonstrates the training loss evolution across four LLM architectures over 80 epochs. Llama2 exhibits superior convergence characteristics, achieving the lowest mean loss by epoch 5 and maintaining consistent improvement throughout training. OPT displays gradual loss reduction with slower optimization dynamics, while T5-LLM follows similar trends but achieves marginally better final loss than OPT. GPT2 initially exhibits the highest loss but rapidly converges from epoch 40, ultimately matching OPT’s performance level.
The loss trajectories validate L4S-LLM’s robust adaptability to complex network queue management across diverse architectural foundations. By leveraging real-time kernel-level metrics, the framework demonstrates effective optimization and generalization under dynamic network conditions. Llama2’s 7B parameter architecture achieves optimal loss minimization, followed by OPT and GPT2 with comparable performance, while T5-LLM exhibits slightly elevated loss values. These results reinforce Llama2’s selection as the primary foundation model, combining superior convergence properties with effective parameter reduction through LoRA for practical deployment in network congestion control scenarios.